# Analysis of the Rankine attraction term in an equation-of-state based on the London dispersion force

### v1

@@ Abstract
The attraction term in an equation of state for gases, $ac^{2}$, proposed by Rankine in 1854, is generally related to the London dispersion force via a calculation of the second virial coefficient, $B_{2}$, by an equation $B_{2}=2\pi N_{0}\int_{0}^{\infty}\left(1-\exp\left(\omega/kT\right)\right)r^{2}\text{d}r$, where $\omega$ is the potential of the attraction between two molecules in the gas. Here we present an alternative approach that does not use this equation, and does not a-priori assume that the function is quadratic in concentration, $c$. Still, the quadratic dependence on concentration is also found. We analyze the gas argon at temperatures between 300 and 800 K. From the numerical calculations, we derive that the attraction parameter depends on temperature according to a -1/6 power scaling, and thus the attraction component to the second virial coefficient, $B_{2}$, scales with a -7/6 power to temperature. For the same conditions, the virial equation presented above results in a square root scaling of $B_{2}$ with temperature, which is less accurate.
@@
In this paper we focus on the attraction term. Rankine himself included that the factor a is inversely proportional to temperature, i.e., it depends on temperature according to $1/T$. Later, Redlich and Kwong (1949) reduced the dependence of a on temperature to a proportionality with $1/\sqrt{T}$, i.e., $a\propto 1/\sqrt{T}$.
The question is whether the quadratic dependence of the attraction term can can be derived from (i.e., is predicted by) the London dispersion force, and what is then the dependence of a on temperature. The London dispersion force describes that the energy of intermolecular attraction depends on one over the sixth power of distance between the molecules. In literature, the expression for the London dispersion force is then implemented in a function $B_{2}=2\pi N_{0}\int\left(1-\exp\left(\omega/kT\right)\right)r^{2}\text{d}r$ ($\omega<0$) to derive the second virial coefficient, $B_{2}$, which apart from a factor RT equals the attraction parameter a. In this approach, it is assumed that the attraction will only be via a quadratic dependence on c. Assuming that $\omega$ is independent of temperature, then for a high enough temperature, $B_{2}$ has an inverse proportionality on T, and thus the virial equation predicts that the attraction parameter, a, is independent of temperature. But for lower temperatures, the predicted dependence of $B_{2}$ and a on temperature can be different.
We evaluate the energy of a gas due to the London dispersion force (attractive force) as follows. Starting from a central molecule, we evaluate the distribution around it of the first, second, third, etc., closest neighbors, and the energy of their interaction, then multiply this result with concentration, to arrive at an energy density (also divide by 2), and take an appropriate derivative to obtain the contribution to pressure. This analysis is very different from the approach discussed above to derive the second virial coefficient via an integration over space of $\exp\left(\omega/kT\right)$.
@@
with $\lambda$ the London dispersion factor, also written as $C\mathrm{{}_{disp}}$, with unit J.m6. It relates to the polarizability of the gas molecules, $\alpha$, and to an energy $h\nu$, according to $\lambda=\tfrac{3}{4}\alpha^{2}h\nu$, where h is the Planck constant and $\nu$ the orbiting frequency of an electron. Dependent on the molecule, this energy is in the range $h\nu=\left\{1.8-3.4\right\}\cdot 10^{-18}$ J. The polarizability of each molecule (atom) is tabulated, such as $\alpha=0.21\penalty 10000\ \AA ^{3}$ for He, and $\alpha=1.64\penalty 10000\ \AA ^{3}$ for Ar. The London dispersion factor, $\lambda$, is independent of temperature. [^1] And for that reason, the Hamaker constant, $A=\pi^{2}q^{2}\lambda$, is temperature-independent as well; here, q is the number density of atoms in a material. In the derivation below, $\lambda$ is equated to the product of kT and a specific length, $r\mathrm{{}_{x}}$, to the power 6, i.e., $\lambda=kT\cdot r_{\text{x}}^{6}$. Because $\lambda$ does not depend on temperature, $r_{\text{x}}^{3}$ is inversely proportional to the square root of temperature, $r_{\text{x}}^{3}\propto 1/\sqrt{T}$.
@@
where the integration limits go from two times the radius of the atoms, $r=2a$ (assuming the atoms are spherical), to infinity. We implement here a function $N_{1}$, which is the nearest neighbor distribution function, given by
@@
where n is a concentration in numbers per volume, $n=c\,N_{0}$, with $N_{0}$ the Loschmidt number ($N_{0}=6.022\cdot 10^{23}$ mol-1). [^2] Eq. (4) is based on a derivation that analyzes the distribution of distances from r=0r\!=\!0 onward. That is different in Eq. (3) because we start at r=2ar\!=\!2a. But we can still use it because the $N_{1}$-function is used in the numerator and denominator in the same way, with the same limits of integration.
@@
Eqs. (5) and (6) only consider interaction with the first neighbor, but we can simply add contributions of the second, third, etc., neighbor. We must only replace in each case the function for ${N}_{1}$ by the general expression for the distribution of the ith neighbor, given by
@@
In a first calculation, we use $\alpha\!=\!1.63\penalty 10000\ \AA ^{3}$ as for Ar, and $h\nu=2.53\cdot 10^{-18}$ J, and thus the London dispersion constant is $\lambda=5.0\cdot 10^{-78}$ J.m6. We calculate $r_{\text{x}}^{3}$ from $r_{\text{x}}^{3}=\sqrt{\lambda/kT}$, and thus at T=300T\!=\!300 K, we have $r_{\text{x}}\!=\!3.27\penalty 10000\ \AA$ . We take a radius $a\!=\!1.72\penalty 10000\ \AA$ , and numerically integrate Eq. (8) from y=0y\!=\!0 (infinitely far) to $y=r_{\text{x}}/\left(2a\right)$, which at 300 K is y=0.950y\!=\!0.950.
@@
We evaluate this attractive pressure due to London dispersion, $P\mathrm{{}_{att}}$, in a range of concentrations, c, from 0 to 500 mM, and find a perfect quadratic dependence, with a prefactor a=0.122a=-0.122 J.m3/mol2. Dividing $a$ by RT results in an attraction contribution to the second virial coefficient of $B\mathrm{{}_{2,att}}=-49$ cm3/mol. That is much more negative than an aggregate value for Ar from literature that is $B_{2}=-16$ cm3/mol, but in that number also the hard sphere effect is included. That contribution we estimate as $B\mathrm{{}_{2,hs}}=34$ cm3/mol, and then $B\mathrm{{}_{2,att,lit}}=-50$ cm3/mol, which is close.
@@
We redo the entire calculation at T=400T\!=\!400 K and 500 K. From the reported aggregate values of $B_{2}$, we again subtract $B\mathrm{{}_{2,hs}}$, and thus assume the literature values for attraction are $B\mathrm{{}_{2,att,lit}}=-35$ and 27-27 cm3/mol, respectively. We again find that the contribution of attraction to the pressure scales with concentration squared, and for $B\mathrm{{}_{2,att}}$ we calculate for 400 and 500 K values of $B\mathrm{{}_{2,att}}=-35$ and 27-27 cm3/mol, respectively, in perfect agreement with the experimental data.
The calculation output shows that the attraction factor a scales with temperature to the power $\tfrac{1}{6}$, i.e., $a\propto 1/\sqrt[6]{T}$, and thus our theoretically derived $B\mathrm{{}_{2,att}}$ scales with temperature to the power $\tfrac{7}{6}$. Also calculations at 600 K and 700 K come out close to the data, with the predicted $B\mathrm{{}_{2,att}}$ given by 22-22 and 18-18 cm3/mol, resp., while the literature data for $B\mathrm{{}_{2,att,lit}}$ are 22-22 and 19-19 cm3/mol, resp. At T=200T\!=\!200 K, we calculate $B\mathrm{{}_{2,att}}=-85$ cm3/mol, while $B\mathrm{{}_{2,att,lit}}=-82$ cm3/mol, again close. (Note that at T=200T\!=\!200 K, the attraction parameter comes out 8% larger than the 1/6th scaling.) At T=160T\!=\!160 K, $B\mathrm{{}_{2,att}}=-122$ cm3/mol, while $B\mathrm{{}_{2,att,lit}}=-110$ cm3/mol, thus a deviation develops now that we come closer to the critical point. (At T=160T\!=\!160 K, the attraction parameter is 20% larger than the 1/6th scaling.) Also of interest is that at these temperatures of 200 K and lower, we need to incorporate more and more ‘neighbors’ to make the pressure curve stay quadratic with concentration. Analyzing the first 4 neighbors as we used at higher temperatures, no longer suffices. Our calculated values for $B_{\text{2,att}}$ are plotted in Fig. 1 as continuous line, together with literature data.
Thus, we propose for Ar in the temperature range 300800300-800 K the equation of state
$P=cRT\cdot\left(1+4\eta+10\eta^{2}+20\eta^{3}\right)-a^{*}\cdot T^{-1/6}\cdot c^{2}$ (10)
@@
$P=cRT\cdot\left(1+4\eta+10\eta^{2}+20\eta^{3}\right)-a^{*}\cdot T^{-1/6}\cdot c^{2}/\left(1+\alpha\eta\right)$ (11)
which gives an improved description of the critical point, especially when we allow the factor $\alpha$ not to be 4, but to be reduced to $\alpha=1.2$. Then the equation predicts $P\mathrm{{}_{c}}=5.5$ MPa, $T\mathrm{{}_{c}}=150$ K, and $c\mathrm{{}_{c}}=13.2$ M, which is very close to the data. Thus, small modifications to the attraction term, that at low concentrations do not have an effect, can markedly modify the prediction of the critical point. In this case, the attraction term is only modified by 10% at the critical point.
Using the virial approach, $B_{2}=2\pi N_{0}\int\left(1-\exp\left(\omega/kT\right)\right)r^{2}\text{d}r$ ($\omega<0$), which we solve using the same atomic radius, a, and same $\lambda$ as above, we obtain $B_{2}=-72$ cm3/mol at 300 K, 61-61 cm3/mol at 400 K, and 54-54 cm3/mol at 500 K. These calculated numbers for $B_{2}$ underestimate the influence of temperature on $B_{2}$, because they scale approximately with $1/\sqrt{T}$. And thus the interaction parameter a is predicted to scale with $\sqrt{T}$, i.e., increases with temperature, which is not so likely.
In conclusion, we presented an analysis of the London dispersion force based on a method that considers interactions of the first, second, third, etc., neighbors around a molecule, and for argon in the temperature range 300800300-800 K, we find a scaling of the attraction parameter, a, in the Rankine attraction term, $ac^{2}$, that is proportional to temperature to the power 1/6-1/6. This result is in agreement with data for $B_{2}$ of argon in that temperature range.
Figure 1: The contribution of attraction to the second virial coefficient of argon, as function of temperature. Continuous line is based on evaluation of the London dispersion force by the method explained in this paper, and circles are literature data. Dashed line is a 7/6th scaling with temperature. These data are obtained from aggregate $B_{2}$-values from which we subtract $B_{\text{2,hs}}=34$ cm3/mol, which is the contribution of hard sphere, volumetric, interactions.

### v2

# Analysis of the Rankine attraction term in an equation of state based on the London dispersion force

@@ Abstract
The attraction term in an equation of state for gases, $ac^{2}$, proposed by Rankine in 1854, is generally related to the London dispersion force via the equation for the second virial coefficient, $B_{2}$, given by $B_{2}=2\pi N_{0}\int_{0}^{\infty}\left(1-\exp\left(-\omega/kT\right)\right)r^{2}\text{d}r$, where $\omega$ is the attraction energy between two molecules in the gas. This equation works very well to describe $B_{2}$ and thus $a=B_{2}RT$ as function of temperature, but the derivation is complicated. Here we present two other methods to derive $B_{2}$ and thus $a$ from the London equation, which have a more intuitive background. (The more simple of the two we expect must be available in literature.) We analyze these three models for the gas argon at temperatures between 150 and 900 K. All three methods fit the data quite closely while an analytical approximation fits data even better. The temperature dependence of $a$ is well described by a proportionality of $a$ with a term $1+T^{*}/T$ where $T^{*}$ is a constant that depends on the type of gas.
@@
In this paper we focus on the attraction term. Rankine assumed that the factor a is inversely proportional to temperature, i.e., it depends on temperature according to $1/T$, i.e., it scales with temperature to the power 1-1. Later, Redlich and Kwong (1949) reduced the dependence of a on temperature to a proportionality with $1/\sqrt{T}$, i.e., $a$ scales with temperature to the power 1/2{-1/2}.
The question is whether the quadratic dependence of the attraction term can can be derived from (i.e., is predicted by) the London dispersion force, and what is then the dependence of a on temperature. The London dispersion force describes that the energy of intermolecular attraction depends on one over the sixth power of distance between the molecules. In literature, the expression for the London dispersion force is then implemented in an equation for the second virial coefficient, $B_{2}$, given by $B_{2}=2\pi N_{0}\int\left(1-\exp\left(-\omega/kT\right)\right)r^{2}\text{d}r$. The attraction part of the second virial coefficient, $B_{\text{2,att}}$, then relates to the attraction parameter a according to $a=B_{\text{2,att}}RT$. In this approach, it is assumed that the attraction term has a quadratic dependence on c. Assuming that $\omega$ is independent of temperature, then for a high enough temperature, $B_{2}$ has an inverse proportionality on T, and thus the virial equation predicts that the attraction parameter, a, is independent of temperature. But for lower temperatures, the predicted dependence of $B_{2}$ and a on temperature can be different.
We will evaluate $B_{2}$ and $a$ of a gas due to the London dispersion force (attractive force) by the virial equation, as well as by two other methods. The advantage of the two alternative methods is that they have a more intuitive background than the virial equation which is based on a quite complicated derivation. The three methods give close to the same results, and give a close fit to data. This correspondence with data, and between the methods, is very important, because it gives support to allowing the use of one method instead of another for other systems of gas molecules or ions.
Of the two alternative methods, we first describe the more detailed and complex method. We consider a gas molecule located at the center of a spherical coordinate system (at r=0r\!=\!0), and we evaluate the distribution around it of the first, second, third, etc., closest neighbors, and calculate the energy of interaction of each of them with the molecule at the center. Then we multiply this result with concentration, to arrive at an energy density (also divide by 2), and take an appropriate derivative to obtain the contribution to pressure.
@@
with $\lambda$ the London dispersion factor, also written as $C\mathrm{{}_{disp}}$, with unit J.m6. It relates to the polarizability of the gas molecules, $\alpha$, and to an energy $h\nu$, according to $\lambda=\tfrac{3}{4}\alpha^{2}h\nu$, where h is the Planck constant and $\nu$ the orbiting frequency of an electron. Dependent on the molecule, this energy is in the range $h\nu=\left\{1.8-3.4\right\}\cdot 10^{-18}$ J. The polarizability of each molecule (atom) is tabulated, such as $\alpha=0.21\penalty 10000\ \AA ^{3}$ for helium, and $\alpha=1.64\penalty 10000\ \AA ^{3}$ for argon. The London dispersion factor, $\lambda$, is independent of temperature. [^1] And for that reason, the Hamaker constant, $A=\pi^{2}q^{2}\lambda$, is temperature-independent as well; here, q is the number density of atoms in a material. In the derivation below, $\lambda$ is equated to the product of kT and a specific length, $r\mathrm{{}_{x}}$, to the power 6, i.e., $\lambda=kT\cdot r_{\text{x}}^{6}$. Because $\lambda$ does not depend on temperature, $r_{\text{x}}^{3}$ is inversely proportional to the square root of temperature, $r_{\text{x}}^{3}\propto 1/\sqrt{T}$.
@@
where the integration limits go from two times the radius of the atoms, $r=2a$ (assuming the atoms are spherical), to infinity. We implement here a function $N_{1}$, which is the nearest neighbor distribution function, given by the Hertz equation
@@
where n is a concentration in numbers per volume, $n=c\,N_{0}$, with $N_{0}$ the Loschmidt number ($N_{0}=6.022\cdot 10^{23}$ mol-1).
@@
Eqs. (5) and (6) only consider interaction with the first neighbor, but we can simply add contributions of the second, third, etc., neighbor. We must only replace in each case the function for ${N}_{1}$ by the general expression for the distribution of the ith neighbor, given by (Berberan Santos, 1987)
@@
In a first calculation, we use $\alpha\!=\!1.63\penalty 10000\ \AA ^{3}$ and $h\nu=2.53\cdot 10^{-18}$ J, which are reported values for argon, and thus the London dispersion constant is $\lambda=5.0\cdot 10^{-78}$ J.m6. We calculate $r_{\text{x}}^{3}$ from $r_{\text{x}}^{3}=\sqrt{\lambda/kT}$, and thus at T=300T\!=\!300 K, we have $r_{\text{x}}\!=\!3.27\penalty 10000\ \AA$ . We take a radius $a\!=\!1.72\penalty 10000\ \AA$ , and numerically integrate Eq. (8) from y=0y\!=\!0 (infinitely far) to $y=r_{\text{x}}/\left(2a\right)$, which at 300 K is y=0.950y\!=\!0.950. (A reported value for the kinetic radius of argon is 0.170 nm, thus we use a value very close to this.)
@@
We make this calculation based on adding together the interaction with the first 5 neighbors (in this temperature range this is enough to have convergence), and evaluate this attractive pressure due to London dispersion, $P\mathrm{{}_{att}}$, in a range of concentrations, c, from 0 to 500 mM, and find a perfect quadratic dependence, with a prefactor a=0.122a=-0.122 J.m3/mol2. Dividing $a$ by RT results in an attraction contribution to the second virial coefficient of $B\mathrm{{}_{2,att}}=-49$ cm3/mol. That is much more negative than an aggregate value for argon from literature that is $B_{2}=-16$ cm3/mol, but in that number also the hard sphere effect is included. That contribution we estimate as $B\mathrm{{}_{2,hs}}=34$ cm3/mol ($B_{\text{2,hs}}=4v\mathrm{{}_{m}}$, with $v\mathrm{{}_{m}}$ a molecular volume), and then $B\mathrm{{}_{2,att,lit}}=-50$ cm3/mol, which is close.
@@
We redo the entire calculation at T=400T\!=\!400 K and 500 K. From the reported aggregate values of $B_{2}$, we again subtract $B\mathrm{{}_{2,hs}}$, and thus assume the literature values for attraction are $B\mathrm{{}_{2,att,lit}}=-35$ and 27-27 cm3/mol, respectively. We again find that the contribution of attraction to the pressure scales with concentration squared, and for $B\mathrm{{}_{2,att}}$ we calculate for 400 and 500 K values of $B\mathrm{{}_{2,att}}=-35$ and 27-27 cm3/mol, respectively, in perfect agreement with the experimental data. Calculations at 600 K and 700 K result in $B\mathrm{{}_{2,att}}$ given by 22-22 and 19-19 cm3/mol, resp., while the literature data for $B\mathrm{{}_{2,att,lit}}$ are exactly the same. At T=200T\!=\!200 K, we calculate $B\mathrm{{}_{2,att}}=-85$ cm3/mol, while $B\mathrm{{}_{2,att,lit}}=-82$ cm3/mol, again close. At T=160T\!=\!160 K, $B\mathrm{{}_{2,att}}=-118$ cm3/mol, while $B\mathrm{{}_{2,att,lit}}=-110$ cm3/mol, thus a deviation develops now that we come closer to the critical point. Interestingly, at these temperatures of 200 K and lower, the quadratic dependence is only closely followed at concentrations up to $\sim 200$ mM, and for the pressure curve to ‘stay quadratic’ up to higher concentrations, we likely have to incorporate more and more neighbors than we did (we always analyzed only the first 5 neighbors).
In an alternative method, we do not consider the first, second, third neighbors, etc., but assume a homogeneous distribution of molecules over space with a background concentration n. The energy of interaction of the central molecule with all molecules around it, $\Psi$, is then given by
$\frac{\Psi}{n}=4\pi\int\exp\left(-\omega/kT\right)\cdot\omega\cdot r^{2}\,\text{d}r$ (10)
where we again integrate from twice the molecular radius, to infinity. To calculate $B_{\text{2,att}}$, we multiply the right-hand side by $N_{0}^{2}$ and divide by $2RT$, which we can rewrite to
$B_{2}=2\pi N_{0}\int\exp\left(-\omega/kT\right)\cdot\left(\omega/kT\right)\cdot r^{2}\,\text{d}r\,.$ (11)
This calculation gives results that are slightly different from the nearest neighbor-method, with $B_{2}$ slightly more negative for $T<300$ K, and slightly less negative for higher temperatures. If we go to temperature of 200 K and lower, it starts to deviate more significantly and gives too negative values of $B_{\text{2,att}}$.
We also solve $B_{\text{2,att}}$ according to the virial approach, $B_{2}=2\pi N_{0}\int\left(1-\exp\left(-\omega/kT\right)\right)r^{2}\text{d}r$, using the same $\lambda$, but a slightly lower molecular radius of $a=0.168$ nm to fit data as good as possible. Data are described very well, only overestimating the data of $B_{\text{2,att}}$ slightly for temperatures below 400 K, up to about 5 point cm3/mol at 120 K. For the homogeneous method and the virial route, the calculated values for $B_{\text{2,att}}$ are plotted in Fig. 1 as a continuous blue line and a red dashed line, respectively, together with literature data. In the full range, the virial route is better, but considering only data above 200 K, the situation is reversed.
We compare in Fig. 2 the argument in the integration of the homogeneous method, Eq. (11), with that of the virial route, $1-\exp\left(\omega/kT\right)$. For low $\omega/kT$, the two methods overlap, but both at lower and higher energies they deviate from one another. At more negative $\omega/kT$, the homogeneous method gives a more negative prediction of $B_{2}$ than the virial route, while for more positive $\omega/kT$, the virial route has an argument that approaches 1, but for the homogeneous route the argument has a maximum and then goes to zero. Because of this difference, the virial route can also correctly predict the hard sphere-repulsion, $B_{\text{2,hs}}=4v\mathrm{{}_{m}}$, but the homogeneous method cannot.
We can make a series expansion of the homogeneous method and the virial route, around $\omega=0$. For low $\omega/kT$, the two methods are the same (the first term is $\omega/kT$), but the second term is $\omega^{2}/\left(kT\right)^{2}$ for the homogeneous route, and half of that for the virial route. After integration of these first two terms for the homogeneous route, we obtain the analytical solution
$B_{\text{2,att}}=-12\cdot v_{\text{m}}^{*}\cdot T^{*}/T\cdot\left(1+T^{*}/T\right)$ (12)
where $v_{\text{m}}^{*}=\frac{4}{3}\pi a^{3}N_{0}$ is a molecular volume based on the radius $a$, and $T^{*}=\lambda/\left(192\cdot a^{6}\cdot k\right)$ is a temperature factor (unit K). When we use $a=0.170$ nm, and thus $v_{\text{m}}^{*}=12.4$ cm3/mol, this function fits the data for argon exactly at all temperatures, better than all methods discussed up to now.
Thus, we propose for argon in the temperature range 150900150-900 K the equation of state
$P=cRT\cdot\left(1+4\,\eta+10\,\eta^{2}+20\,\eta^{3}\right)-12\cdot R\cdot v_{\text{m}}^{*}\cdot T^{*}\cdot\left(1+T^{*}/T\right)\cdot c^{2}$ (13)
where $\eta=v\mathrm{{}_{m}}c$ is based on $v\mathrm{{}_{m}}=8.5$ cm3/mol, while for argon, $T^{*}=79.6$ K. Here we also added the next two terms in the Jäger expression for the hard sphere repulsion (nowadays, the factor 20 is replaced by 18 in the Carnahan-Starling equation of state). With these numbers, argon is athermal at $T=415$ K, i.e., in the dilute limit, volume effects and attraction cancel each other out at that temperature, and it then behaves as an ideal gas. Based on Eq. (13), the critical point is when $dP/dc=0$ and $d^{2}P/dc^{2}=0$, which results in $P\mathrm{{}_{c}}=8.9$ MPa, $T\mathrm{{}_{c}}=187$ K, and $c\mathrm{{}_{c}}=16.4$ M. Experimental data are $P\mathrm{{}_{c,exp}}=4.9$ MPa, $T\mathrm{{}_{c,exp}}=150.8$ K, and $c\mathrm{{}_{c,exp}}=13.4$ M, and thus this simple theory that was only evaluated at low concentrations (<0.5<0.5 M) is definitely off, but not dramatically.
@@
which gives an improved description of the critical point, especially when we allow the factor $\alpha$ not to be 4, but to be reduced to $\alpha\!=\!1.9$. Then the equation predicts $P\mathrm{{}_{c}}=4.9$ MPa, $T\mathrm{{}_{c}}=149$ K, and $c\mathrm{{}_{c}}=12.0$ M, which are predictions close to the data. Thus, small modifications to the attraction term, that at low concentrations do not have an effect, can markedly modify the prediction of the critical point. In this case, the attraction term is modified by $\sim 15\%$ at the critical point when we include $1+\alpha\eta$. The temperature factor $T^{*}$ depends on the type of gas via the kinetic radius, $a$, and London coefficient, $\lambda$. We calculate for helium, neon, argon, and xenon, the following temperature factors: $T^{*}=15.7,\penalty 10000\ 22.0,\penalty 10000\ 79.6,\penalty 10000\ \text{and}\penalty 10000\ 128.3$ K, resp.
In conclusion, we presented an analysis of the London dispersion force based on the virial route commonly used in literature, and two alternative methods, one where a homogeneous concentration is considered, and one including in more detail interactions of a molecule with its first, second, third, etc., neighbors. For argon in the temperature range 150900150-900 K, we find that all methods are close to the data, while an analytical equation fits data even better. In this analytical equation, the attraction parameter, a, in the Rankine attraction term, $ac^{2}$, depends on temperature according to $a\propto 1+T^{*}/T$ with T a material-specific temperature factor. This dependence implies that the scaling of a with temperature is for instance $a\propto T^{-1/10}$ at a temperature of 700 K, and $a\propto T^{-1/5}$ at 300 K. A small volume correction to the attraction term due to Redlich-Kwong has a marked effect on the prediction of the critical point of argon.
Figure 1: The contribution of attraction to the second virial coefficient of argon, as function of temperature. The red dashed line is according to the virial route, and continuous blue line is based on evaluation of the London dispersion force by the homogeneous method. Circles are data from literature of aggregate $B_{2}$-values from which a hard sphere-contribution, $B_{\text{2,hs}}=34$ cm3/mol, is subtracted.
Figure 2: Comparison of the virial route (dashed red line) with the homogeneous method, Eq. (11) (continuous blue line), plotting the argument in the integration as function of $\omega/kT$ (ignoring $r^{2}$). The two methods overlap for small $\omega/kT$, but they are different for lower and for higher energies.